Circular No. 7712

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001eg IN UGC 3885 G. M. Hurst, Basingstoke, England, reports the discovery by M. Arm-

strong (Rolvenden, England) of an apparent supernova (mag 17.8) on an unfiltered CCD image (limiting mag 19) taken on Aug. 29.138 UT in the course of the U.K. Nova/Supernova Patrol with a 0.36-m Schmidt-Cassegrain telescope. The discovery image yields the following position for the new object: α = 7^{h}31^{m}05^{s}.91, δ = +59^{o}28′38′′.8 (equinox 2000.0), which is 7′′ west and 13′′ south of the center of UGC 3885. A CCD image taken by Armstrong on 2000 Dec. 21 shows no object at this location to limiting mag 19.5. An image taken by T. Boles (Coddenham, England) on Sept. 1.05 seems to show the new object, from which Armstrong provides the following position end figures: 05^{s}.95, 38′′.7. P. Meikle, Astrophysics Group, Imperial College of Science, London, reports that two 600-s I-band images (limiting mag 21.1) obtained by D. Pollacco with the 1.0-m Jacobus Kapteyn Telescope at La Palma on Sept. 2.2 UT show SN 2001eg as a point source at mag 19.2 ± 0.3 and located 6′′.8 west and 16′′.0 south of the galaxy nucleus; Armstrong has measured the position from these images and provides the following position end figures: 05^{s}.99, 38′′.3. Armstrong adds that an image that he obtained on Sept. 5.14 shows the new object at mag 18.7.

SUPERNOVA 2001eh IN UGC 1162 Hurst also reports the discovery by Armstrong of another apparent

supernova (mag 16.5) on an unfiltered CCD image (limiting mag 18.5, affected by slight trailing) taken on Sept. 9.119 UT as above with a 0.30-m Schmidt-Cassegrain telescope; the new object appeared at the same brightness on another image secured on Sept. 10.050. The following position for SN 2001eh is the average of measures by Armstrong from his two images: α = 1^{h}38^{m}12^{s}.10, δ = +41^{o}39′18′′.8 (equinox 2000.0), which is 36′′ west and 6′′ north of the center of UGC 1162. Nothing appears at this position on an image taken by Armstrong on 2000 Sept. 25 (limiting mag 19.5), and nothing is anything visible on Palomar Sky survey images taken on 1991 Jan. 19 (blue, limiting mag 22.5), 1995 May 19 (infrared), and 1998 Nov. 25 (red, limiting mag 20.8).

V2275 CYGNI CCD photometry by G. Sostero and A. Lepardo, Udine, Italy: Sept.

5.80 UT, V = 10.53, B − V = +0.9, V − R_{c} = +1.1, R_{c} − I_{c} = +0.9; 6.87, 10.60, +0.9, +1.0, +1.0; 9.82, 10.83, +0.9, +1.1, +0.8.

2001 September 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green